1/10(x+8) - 1/15(x-5)=1 My teacher told me the answer was -4 I think...I got negative 33 which is obviously wrong xD. Please help me know my error and see how to get the right answer. I'm usually good with this equations, but it was probably a simple mistake that gave me my subsequent error. Okay, well, if you answer, I would be most delightful and thankful!!
Answer by stanbon(75887)

(1/10)(x+8) - (1/15)(x-5)=1 My teacher told me the answer was -4
-------------
Multiply thru by 30 to get rid of the fractions:
3(x+8) - 2(x-5) = 30
3x + 24 - 2x + 10 = 30
x + 34 = 30
x = -4
===========
Cheers,
Stan H.